# On Probabilistic Pullback Metrics on Latent Hyperbolic Manifolds

### v1

@@ 4 Hyperbolic Kernel Derivatives
where $\bm{G}=\text{diag}(-1,1,...,1)$ is the Lorentzian metric tensor. The helper functions $g(u)=\left(\frac{2\rho^{2}}{\kappa s^{2}}-\frac{1}{s^{2}}-\frac{u\rho}{s^{3}}\right)$, and $h(u)=\frac{\text{d}}{\text{d}u}g(u)+\frac{g(u)2\rho}{\kappa s}$ are essential to understand why standard autodiff tools are unable to compute these derivatives. For equal kernel inputs $\bm{x}=\bm{z}$ the Lorentzian inner product $u$ approaches $-1$ and both the distance $\rho$ and helper variable $s$ converge to 00. Analytically for the helper functions this limit is well behaved, e.g., $\lim_{u\rightarrow-1^{-}}g(u)=\frac{2}{\kappa}+\frac{1}{3}$. However, autodiff tools like PyTorch fail to compute these derivatives correctly because they are unable to cancel out the 00-approaching terms. Take for example $g(u)$,
$\dfrac{2\rho^{2}}{\kappa s^{2}}-\dfrac{1}{s^{2}}-\dfrac{u\rho}{s^{3}}\rightarrow^{\text{ autodiff}}_{\>\>\>\>\bm{x}=\bm{z}}\>\rightarrow\dfrac{0}{0}-\dfrac{1}{0}-\dfrac{0}{0}=\text{NaN}\,.$
@@
In this scenario, autodiff ends up dividing by zero, resulting in undefined values (NaN). In other cases, e.g., when training a GPLVM, derivatives for equal kernel inputs $\bm{x}=\bm{z}$ may not be that relevant because different latent points rarely become that close. However, these derivatives are essential for computing the Jacobian covariance matrix (18). Although using symbolic differentiation libraries could provide a solution, they tend to be too slow for practical purposes. Therefore, the next best approach is to compute the necessary derivatives and their analytical limits manually. The analytical limits are then used whenever the inputs $\bm{x},\bm{z}$ are closer than a predefined threshold.
@@
Finally, in order to optimize geodesics on the pullback metric we need the third-order kernel derivatives like in the 2D case. Let $\bm{K}=\frac{\partial^{2}}{\partial\bm{z}\partial\bm{x}}k(\bm{x},\bm{z})\in\mathbb{R}^{4\times 4}$ now be the second kernel derivative. The third-order derivatives can then be obtained by stacking the individual derivatives of each row and column of $\bm{K}$. The derivative of each row $\bm{K}_{i}$ w.r.t. the first input $\bm{x}$ and each column $\bm{K}_{j}$ w.r.t the second input $\bm{z}$ is given by,
$\!\!\dfrac{\partial\bm{K}_{i}}{\partial\bm{x}}\!=\!\Big{(}\!q(\!u\!)\bm{G}\bm{z}\bm{x}^{\!\mathsf{T}}\!\bm{G}z_{i}\!+\!h(\!u\!)\bm{G}z_{i}\!+\!h(\!u\!)\bm{Gz}\bm{e}_{i}^{\!\mathsf{T}}\!\Big{)}\,\!\Delta_{1i}\!\,e^{\frac{-\rho^{2}}{\kappa}}\!,$
$\!\!\dfrac{\partial\bm{K}_{\!j}}{\partial\bm{z}}\!=\!\Big{(}\!q(\!u\!)\bm{G}\bm{z}\bm{x}^{\!\mathsf{T}}\!\bm{G}x_{\!j}\!+\!h(\!u\!)\bm{G}x_{\!j}\!+\!h(\!u\!)\bm{e}_{\!j}\bm{x}^{\!\mathsf{T}}\!\bm{G}\!\Big{)}\Delta_{1j}e^{\frac{-\rho^{2}}{\kappa}}\!,$
@@
where $q(u)=\frac{\text{d}}{\text{d}u}h(u)+\frac{h(u)2\rho}{\kappa s}$ and $\Delta_{ij}$ represents a modified Kronecker delta equal to $-1$ for equal indices and to $1$ otherwise. Stacking the row and column matrices $\bm{K}_{i}$ and $\bm{K}_{j}$ along the first and second dimension accordingly gives the third kernel derivatives $\frac{\partial}{\partial\bm{x}}\bm{K},\frac{\partial}{\partial\bm{z}}\bm{K}\in\mathbb{R}^{4\times 4\times 4}$.
@@
Figure 2: Left: 2D hyperbolic latent space of a subset of the MNIST dataset with $\mathsf{0}$ ( ), $\mathsf{1}$ ( ), $\mathsf{2}$ ( ), $\mathsf{3}$ ( ), $\mathsf{6}$ ( ), and $\mathsf{9}$ ( ). Both the hyperbolic ( ) and pullback ( ) geodesics interpolate between a $\mathsf{3}$ and a $\mathsf{6}$. Middle: Zoomed-in view of the latent space to better visualize the metric volume. Right: Ten samples along the decoded geodesics in the image space, corresponding to the hyperbolic geodesic (top rows) and the the pullback geodesic (bottom rows).
@@ 5 Experiments
In this section we present experiments to demonstrate that (1) The hyperbolic pullback metric augments the hyperbolic metric with the distortions introduced by the GPHLVM’s nonlinear mapping; and (2) Similar to Euclidean pullback metrics, the hyperbolic pullback metric generates geodesics that adhere to the training data, allowing low uncertainty model predictions. Further experimental details are provided in App. C.
@@ 5.1 C-shape
The first experiment serves as a proof of concept to visualize and compare the Euclidean and hyperbolic pullback metrics from a GPLVM and GPHLVM, respectively. We use a dataset of $N$ two-dimensional $\mathsf{C}$-shape points as both latent points and observations. In the hyperbolic case, the latent points and observations are represented in the Lorentz model, i.e, $\bm{x},\bm{y}\in\mathbb{H}^{2}_{\mathcal{L}}$. Only for visualization purposes we transform the latent points into the Poincaré model. To fully specify the models, we set the values for the variance, length scale, and noise variance, as $(\tau,\kappa,\sigma_{y}^{2})=(0.7,0.15,0.69)$.
Figure 1 displays the Euclidean and hyperbolic pullback metric volumes. To generate these plots, we sampled a grid of points within the unit circle. Each grid point serves as a latent point $\bm{x}^{*}$, for which we compute the pullback metric tensor $\bm{G}_{\bm{x}^{*}}$ via Equation (20). The volume of $\bm{G}_{\bm{x}^{*}}$ is then given by $\sqrt{\text{det}(\bm{G}_{\bm{x}^{*}})}$. In the hyperbolic case, the pullback metric tensor is a $3\times 3$ matrix. Since this matrix lies in the two-dimensional tangent space of $\bm{x}^{*}$, one of its eigenvalues is always zero. To effectively visualize the volume, we exclude the zero eigenvalue. We observe in Figure 1 that, in the Euclidean case, the pullback metric volume is small near the C-shape data and increases away from it, until a maximum value is reached. In the hyperbolic case, the metric volume is also low nearby the data. However, it additionally follows the hyperbolic geometry: It is low near the origin and increases when moving outwards until it becomes infinite at the boundary of the unit circle. Overall, the hyperbolic pullback metric effectively integrates the properties of the hyperbolic manifold with those of the data manifold.
Figure 1 also depicts geodesics on both latent manifolds, which are generated with and without the pullback metric (solid and dashed black lines, respectively). In both cases, the pullback geodesic closely adheres to the training data. However, in the hyperbolic case, crossing the center experiences a much lower penalty, in terms of curve energy, than in the Euclidean case. The rightmost plots in Fig. 1 display the curve energy (21) along the geodesics. We observe that the geodesics on the base manifold and the pullback geodesic (dashed and solid black curves, respectively) both exhibit constant energy, confirming that the optimized curves are true geodesics on their respective manifolds. For comparison, we also show the curve energy of the base manifold geodesic, but evaluated using the pullback metric tensor (dashed green curve). Its energy is clearly not constant, indicating that the base manifold geodesic is not a true geodesic with respect to the pullback metric.
@@ 5.2 MNIST Digits
The natural clustering of MNIST images can be viewed as a hierarchy whose nodes are each of the $10$ digit classes. From this point of view, hyperbolic spaces are well suited for embedding MNIST images (Mathieu et al.,, 2019). Similarly to previous works (Arvanitidis et al.,, 2018; Jørgensen and Hauberg,, 2021; Lalchand et al.,, 2022), we here explore the interpolation of handwritten digits from a subset of the MNIST dataset. Each $28\times 28$ image is vectorized and embedded via the GPHLVM. Figure 2 shows the 2D hyperbolic latent space with two geodesics interpolating between a sample of digit $\mathsf{3}$ and a sample of digit $\mathsf{6}$. The decoded hyperbolic geodesic (dashed black) and pullback geodesic (black), sampled at $10$ time steps, show the result of geodesic interpolation in the image space.
As expected, the GPHLVM predictions collapse to the non-informative GP mean in regions with sparse data, resulting in blurry images. This is particularly evident along the hyperbolic geodesic, in the first half of the trajectory $(t=0.11\text{ to }t=0.44)$. In contrast, the pullback geodesic tends to avoid empty regions by bending towards the red cluster (digit $\mathsf{9}$), leading to less blurry predictions overall. However, since MNIST mainly consists of distinct clusters without smooth transitions, the interpolation setting might be ill-posed. Despite this, our metric better captures the data manifold geometry, and it could thus be used to, for example, measure manifold-aware distances between latent variables.
@@ 5.3 Multi-cellular Robot Design
Figure 3: Top: 2D hyperbolic latent space of multicellular robot embeddings from coarse (darker tone) to fine (lighter tone). The embeddings form two clusters originating from an all-vertically-actuated robot ( ) and an all-horizontally-actuated robot ( ). Bottom: Ten samples along the decoded hyperbolic ( ) and pullback ( ) geodesic.
We build on the work of Dong et al., (2024), which introduced a coarse-to-fine framework for designing multicellular robots using hyperbolic embeddings. Each robot is composed of a $5\times 5$ grid of cells, where each cell can be horizontally actuated (orange), vertically actuated (blue), rigid (black), soft (gray), or empty (white). The design process begins with a fully horizontally or vertically actuated robot and incrementally introduces changes, forming a hierarchical structure that refines from coarse to fine architectures. This hierarchy can then be queried to find suitable robots for specific tasks.
The main goal of this experiment is to leverage pullback geodesics as a data augmentation mechanism, so that we can design new robot architectures by decoding geodesics that follow the pattern of previously-designed robots. To do so, we compute the pullback metric from the hierarchical multicellular robot embeddings and optimize geodesics to interpolate between existing robot designs, generating novel ones. Figure 3 shows the latent space and the mean predictions along both the hyperbolic and pullback geodesics. Note that, while the hyperbolic geodesic produces invalid robot designs, e.g., separated sections at $t=0.22$, the pullback geodesic avoids this issue by complying with the geometry of the data support. This experiment suggests that pullback geodesics can assist the robot design process.
@@ 5.4 Hand Grasps Generation
Figure 4: Hand grasps transition using geodesics optimized on the pullback metric ( ) and the hyperbolic geodesic ( ), from a ring grasp ( ) to a spherical grasp ( ). Top: 2D and 3D latent spaces of the trained models, where the latent grasp trajectory points are coloured according to their corresponding grasp class. In the 2D plot, the background color represents the pullback metric volume. Bottom: Time-series plots of $3$ dimensions of the joint space. The top and bottom rows show the mean of the decoded hyperbolic and pullback geodesics ( and ) and their uncertainty as a gray envelope. For reference, a training trajectory to the spherical grasp ( ) and a reversed training trajectory from the ring grasp ( ) are included.
Our final experiment centers around the grasps taxonomy of Stival et al., (2019), which categorizes human hand grasps into $20$ distinct classes, organized within a hierarchical taxonomy graph. These grasps classes are further grouped into five categories: Flat grasps, distal grasps, cylindrical grasps, spherical grasps, and ring grasps. For each of the $20$ grasp types, the hand movements of multiple subjects were recorded and made available through the KIT whole-body motion database (Mandery et al.,, 2016). For each grasp, subjects started the motion from the same resting pose before reaching out to grasp an object. The recorded motion resulted into trajectories $\bm{Y}\in\mathbb{R}^{N\times 24}$ in a $24$-dimensional space, corresponding to the $24$ degrees of freedom of the hand model. Training two GPHLVMs with dynamic prior akin to (Wang et al.,, 2008) on this dataset results in the 2D and 3D latent spaces shown in Fig. 4. All trajectories in the latent space start near the origin, representing the hand’s initial resting pose, and progress outward. The final point of each trajectory reflects the final grasp.
The goal of this experiment is to use geodesics as a motion generation mechanism, thus enabling new motions that transition from one grasp to another. Figure 4 shows two geodesics interpolating from a ring grasp to a spherical grasp. While the hyperbolic geodesic crosses empty (high uncertainty) space, the pullback geodesic adheres more closely to the data support. We then decode the latent geodesics into the $24$-dimensional hand joint space, to analyze the resulting hand motion. For a full analysis of all $24$ predicted dimensions we refer to App. C.
@@
As both geodesics start and end at the same grasps, their predictions match at the beginning and the end. However, in the middle, the decoded hyperbolic geodesic shows high uncertainty. This is due to the GP reverting to its non-informative mean in regions with no data, where the GP prior dominates the posterior, leading to maximum uncertainty. In contrast, the decoded pullback geodesic shows much lower uncertainty, as the geodesic stays closer to the data support. Despite the low uncertainty, the predicted motions are not particularly smooth. They sometimes exhibit jerky motion trajectories produced by the fact that nearby latent points do not always correspond to equally similar hand configurations (i.e., similar joint values). However, the issue is partially mitigated in the 3D latent space, which provides more room to accommodate the latent embeddings. Improving the smoothness of trajectory predictions still remains an open problem for future work.

### v2

@@ 4 Hyperbolic Kernel Derivatives
with $g(u)\!=\!\left(\frac{2\rho^{2}}{\nu s^{2}}-\frac{1}{s^{2}}-\frac{u\rho}{s^{3}}\right)$ and $h(u)\!=\!\frac{\text{d}}{\text{d}u}g(u)+\frac{g(u)2\rho}{\nu s}$. The helper functions $g(u)$ and $h(u)$ are essential to understand why standard autodiff tools fail to compute the derivatives (29)-(30). For equal inputs $\bm{x}\!=\!\bm{z}$, the inner product $u$ approaches $-1$, while the distance $\rho=\text{d}_{\mathbb{H}^{D_{x}}_{\mathcal{L}}}(\bm{x},\bm{z})$ and variable $s$ converge to 00. This limit is analytically well behaved for the helper functions, e.g., $\lim_{u\rightarrow-1^{-}}g(u)=\frac{2}{\nu}+\frac{1}{3}$. However, autodiff tools fail to compute the kernel derivatives for equal inputs correctly because they do not cancel out 00-approaching terms. For example, we have for $g(u)$,
$\dfrac{2\rho^{2}}{\nu s^{2}}-\dfrac{1}{s^{2}}-\dfrac{u\rho}{s^{3}}\rightarrow^{\text{ autodiff}}_{\>\>\>\>\bm{x}=\bm{z}}\>\rightarrow\dfrac{0}{0}-\dfrac{1}{0}-\dfrac{0}{0}=\text{NaN}\,.$
@@
In this scenario, autodiff divides by 00, leading to undefined values (NaN). In other cases, e.g., when training a GPLVM, derivatives for equal kernel inputs $\bm{x}=\bm{z}$ may not be that relevant because different latent points rarely become that close. However, these derivatives are essential for computing the Jacobian covariance matrix (20). Although symbolic differentiation libraries could provide a solution, they tend to be too slow for practical purposes. Therefore, the next best approach is to compute the necessary derivatives and their analytical limits manually. The latter are then used when the inputs $\bm{x},\bm{z}$ are closer than a predefined threshold.
@@
Finally, we need the third-order kernel derivatives to optimize pullback geodesics as in the 2D case. We denote the second kernel derivative as $\bm{K}=\frac{\partial^{2}}{\partial\bm{z}\partial\bm{x}}k(\bm{x},\bm{z})\in\mathbb{R}^{4\times 4}$. The third-order derivatives can then be obtained by stacking the individual derivatives of each row and column of $\bm{K}$. The derivative of each row $\bm{K}_{i}$ w.r.t. the first input $\bm{x}$ and each column $\bm{K}_{j}$ w.r.t the second input $\bm{z}$ are given by,
$\!\!\dfrac{\partial\bm{K}_{i}}{\partial\bm{x}}\!=\!\Big{(}\!q(\!u\!)\bm{G}^{\mathcal{L}}\bm{z}\bm{x}^{\!\mathsf{T}}\!\bm{G}^{\mathcal{L}}z_{i}\!+\!h(\!u\!)\bm{G}^{\mathcal{L}}z_{i}\!+\!h(\!u\!)\bm{G}^{\mathcal{L}}\bm{z}\bm{e}_{i}^{\!\mathsf{T}}\!\Big{)}\,\!\Delta_{1i}\!\,e^{\frac{-\rho^{2}}{\nu}}\!,$
$\!\!\dfrac{\partial\bm{K}_{\!j}}{\partial\bm{z}}\!=\!\Big{(}\!q(\!u\!)\bm{G}^{\mathcal{L}}\bm{z}\bm{x}^{\!\mathsf{T}}\!\bm{G}^{\mathcal{L}}x_{\!j}\!+\!h(\!u\!)\bm{G}^{\mathcal{L}}x_{\!j}\!+\!h(\!u\!)\bm{e}_{\!j}\bm{x}^{\!\mathsf{T}}\!\bm{G}^{\mathcal{L}}\!\Big{)}\Delta_{1j}e^{\frac{-\rho^{2}}{\nu}}\!,$
@@
where $q(u)=\frac{\text{d}}{\text{d}u}h(u)+\frac{h(u)2\rho}{\nu s}$ and $\Delta_{ij}$ represents a modified Kronecker delta equal to $-1$ for equal indices and to $1$ otherwise. Stacking the row and column matrices $\bm{K}_{i}$ and $\bm{K}_{j}$ along the first and second dimension accordingly gives the third kernel derivatives $\frac{\partial}{\partial\bm{x}}\bm{K},\frac{\partial}{\partial\bm{z}}\bm{K}\in\mathbb{R}^{4\times 4\times 4}$. Complete derivations are provided in App. B.
@@
Figure 3: Left: Latent embeddings in $\mathbb{H}^{2}_{\mathcal{P}}$ of a subset of the MNIST dataset with $\mathsf{0}$ ( ), $\mathsf{1}$ ( ), $\mathsf{2}$ ( ), $\mathsf{3}$ ( ), $\mathsf{6}$ ( ), and $\mathsf{9}$ ( ). The hyperbolic ( ) and pullback ( ) geodesics interpolate between a $\mathsf{3}$ and a $\mathsf{6}$. Middle: Zoomed-in view to better visualize the metric volume. Right: Ten samples along the decoded geodesics in the image space, corresponding to the hyperbolic (top) and pullback (bottom) geodesic.
@@ 5 Experiments
We test the proposed pullback metric in four distinct experiments to demonstrate that (1) the hyperbolic pullback metric augments the hyperbolic metric with the distortions introduced by the GPHLVM’s nonlinear mapping; and (2) the hyperbolic pullback geodesics adhere to the data distribution, leading to low uncertainty model predictions. Additional experimental details are provided in App. D.
@@ 5.1 C-shape
The first experiment serves as a proof of concept to visualize and compare the Euclidean pullback metric $\bm{G}^{\text{P},\mathbb{R}}$ from a GPLVM and the hyperbolic pullback metric $\bm{G}^{\text{P},\mathcal{L}}$ from a GPHLVM. We use a dataset of $N$ 2D $\mathsf{C}$-shape points as both latent variables and observations. In the hyperbolic case, the latent variables and observations are encoded in the Lorentz model, i.e, $\bm{x},\bm{y}\in\mathbb{H}^{2}_{\mathcal{L}}$, but visualized in the Poincaré model which provides a more intuitive representation. We fully specify the LVMs by setting the variance, length scale, and noise variance as $(\tau,\kappa,\sigma_{y}^{2})=(0.7,0.15,0.69)$.
Fig. 2 displays the Euclidean and hyperbolic pullback metric volumes $\sqrt{\text{det}(\bm{G}^{\text{P},\mathbb{R}})}$ and $\sqrt{\text{det}(\bm{G}^{\text{P},\mathcal{L}})}$. In the hyperbolic case, the pullback metric is a $3\times 3$ matrix lying in the 2D tangent space $\mathcal{T}_{\bm{x}^{*}}\mathbb{H}^{2}_{\mathcal{L}}$. Therefore, one of its eigenvalues is always zero, which we exclude to effectively visualize the volume. As shown in Fig. 2-left, the Euclidean pullback metric volume is small near the C-shape data and increases away from it, until a maximum value is reached. The hyperbolic pullback metric volume is also low nearby the data and additionally follows the hyperbolic geometry: It is low near the origin and increases when moving outwards until it becomes infinite at the boundary of the unit circle (see Fig. 2-middle). Overall, the hyperbolic pullback metric effectively integrates the properties of the hyperbolic manifold with those of the data manifold.
Fig. 2 also depicts geodesics generated with and without the pullback metric on both latent manifolds. In both cases, the pullback geodesic closely adheres to the training data. However, in the hyperbolic case, crossing the center experiences a much lower penalty, in terms of curve energy, than in the Euclidean case. Fig. 2-right display the curve energy (23) along the geodesics. We observe that the base manifold geodesics and the pullback geodesics both exhibit constant energy, confirming that the optimized curves are true geodesics on their respective manifolds. For comparison, we also show the curve energy of the base manifold geodesic, but evaluated using the pullback metric tensor. Its varying energy indicates that the base manifold geodesic is not a true geodesic with respect to the pullback metric.
@@ 5.2 MNIST Digits
The natural clustering of MNIST images can be viewed as a hierarchy whose nodes are each of the $10$ digit classes. Following this point of view, Mathieu et al. (2019) showed that hyperbolic LVMs are better suited than their Euclidean counterparts for embedding MNIST images. Similarly to previous works (Arvanitidis et al., 2018; Jørgensen & Hauberg, 2021; Lalchand et al., 2022), we here explore the interpolation of handwritten digits from a subset of the MNIST dataset. We embedded the vectorized $28\times 28$ images in $\mathbb{H}^{2}_{\mathcal{L}}$ via a GPHLVM and then use the induced pullback metric.
Fig. 3 shows the learned hyperbolic latent space along with two geodesics interpolating between a digit $\mathsf{3}$ and a digit $\mathsf{6}$. The decoded hyperbolic geodesic and pullback geodesic, sampled at $10$ time steps, show the result of geodesic interpolation in the image space. As expected, the GPHLVM predictions collapse to the non-informative GP mean in regions with sparse data, resulting in blurry images. This is particularly evident along the hyperbolic geodesic, in the first half of the trajectory $(t=[0.11,0.44])$. In contrast, the pullback geodesic tends to avoid empty regions by bending towards the red cluster (digit $\mathsf{9}$), leading to less blurry predictions overall. This shows that the pullback metric better captures the data manifold geometry, and thus may be used, e.g., to measure manifold-aware latent distances and probabilities.
@@ 5.3 Multi-cellular Robot Design
Figure 4: Top: Embeddings of multicellular robots from coarse (darker tone) to fine (lighter tone) in $\mathbb{H}^{2}_{\mathcal{P}}$. The embeddings form two clusters originating from an all-vertically-actuated robot ( ) and an all-horizontally-actuated robot ( ). Bottom: Ten samples along the decoded hyperbolic ( ) and pullback ( ) geodesic.
We build the coarse-to-fine framework for designing multicellular robots using hyperbolic embeddings introduced by Dong et al. (2024). Each robot is composed of a $5\times 5$ grid of cells, where each cell can be horizontally actuated (orange), vertically actuated (blue), rigid (black), soft (gray), or empty (white). The design process begins with a fully horizontally or vertically actuated robot and incrementally introduces changes, forming a hierarchical structure that refines from coarse to fine architectures. This hierarchy can then be queried to find suitable robots for specific tasks. Dong et al. (2024) showed that their multicellular robot design approach in hyperbolic space outperformed its Euclidean counterpart in most of the considered design tasks.
The main goal of this experiment is to leverage pullback geodesics as a data augmentation mechanism, so that we can design new robot architectures by decoding geodesics that follow the pattern of previously-designed robots. To do so, we compute the pullback metric from the hierarchical multicellular robot embeddings and optimize geodesics to interpolate between existing robot designs, generating novel ones. Fig. 4 shows the latent space and the mean predictions along both the hyperbolic and pullback geodesics. Note that, while the hyperbolic geodesic produces invalid robot designs, e.g., separated sections at $t=0.22$, the pullback geodesic avoids this issue by complying with the geometry of the data support. This experiment suggests that pullback geodesics can assist the robot design process.
@@ 5.4 Hand Grasps Generation
(a) GPHLVM, $\mathbb{H}^{2}_{\mathcal{L}}$
(b) GPHLVM, $\mathbb{H}^{3}_{\mathcal{L}}$
@@
Figure 5: Hand grasps transition from a ring ( ) to a spherical grasp ( ) using hyperbolic ( ) and pullback ( ) geodesics in (a) $\mathbb{H}^{2}_{\mathcal{L}}$ and (b) $\mathbb{H}^{3}_{\mathcal{L}}$. Top left: Latent embeddings. The latent trajectory points are colored according to their corresponding grasp class. The background color in $\mathbb{H}^{2}_{\mathcal{P}}$ represents the pullback metric volume. Top right: Time-series plots of $2$ dimensions of the joint space. The top and bottom rows show the mean of the decoded hyperbolic ( ) and pullback geodesics ( ) with their uncertainty as a gray envelope. A training trajectory to the spherical grasp ( ) and a reversed training trajectory from the ring grasp ( ) are included for reference. Bottom: Generated hand trajectories from the decoded hyperbolic (top) and pullback (bottom) geodesics.
Jaquier et al. (2024) showed that GPHLVM outperform Euclidean GPLVM to learn hand grasp embeddings that comply with the hierarchical structure of human motion taxonomies. Here, we build on this model and explore the use of pullback geodesics as a motion generation mechanism with the aim of creating new motions that transition from one grasp to another. We consider the hand grasp taxonomy of Stival et al. (2019), which organizes common grasps types within a hierarchical taxonomy tree. We use a dataset from the KIT whole-body motion database (Mandery et al., 2016) consisting of $38$ motions of $19$ grasp types obtained from recordings of humans grasping different objects. Each motion corresponds to a subject reaching out to grasp an object from an initial resting pose and results in a trajectory $\bm{Y}\in\mathbb{R}^{N\times 24}$ representing the temporal evolution of the $24$ degrees of freedom of the wrist and fingers.
We train two GPHLVMs augmented with a taxonomy graph-distance prior (Jaquier et al., 2024) and a dynamic prior akin to (Wang et al., 2008) (see App. D.4). Fig. 5 shows the resulting 2D and 3D latent spaces. All training trajectories start near the latent space origin, corresponding to the hand’s initial resting pose, and progress outward until the final grasp. We evaluate the embedding quality via the stress loss $\ell_{\text{stress}}$, that quantifies the preservation of the taxonomy graph distances by the embeddings (see App. D.4 for details). Next, we compare the GPHLVMs with GPLVMs (see also App. D.5). Table 2 shows that the hyperbolic models better capture the taxonomy structure than their Euclidean counterparts, consistent with (Jaquier et al., 2024).
Table 2: Average stress and prediction uncertainty (along decoded geodesics). The stress is compared across geometries for each latent space dimension. The prediction uncertainty is compared between the base manifold and pullback geodesic for each model.
Model Stress $\downarrow$ Prediction Uncertainty $\downarrow$
Base Manifold Pullback
GPLVM $\mathbb{R}^{2}$ $0.18\pm 0.47$ $0.0524\pm 0.0199$ $\bm{0.0004\pm 0.0002}$
GPHLVM $\mathbb{H}^{2}_{\mathcal{L}}$ $\bm{0.10\pm 0.26}$ $0.0150\pm 0.0227$ $\bm{0.0004\pm 0.0002}$
GPLVM $\mathbb{R}^{3}$ $0.11\pm 0.21$ $0.0030\pm 0.0036$ $\bm{0.0001\pm 0.0001}$
GPHLVM $\mathbb{H}^{3}_{\mathcal{L}}$ $\bm{0.07\pm 0.10}$ $0.0129\pm 0.0150$ $\bm{0.0003\pm 0.0003}$
@@
Fig. 5 shows the hyperbolic and pullback geodesics interpolating from a ring to a spherical grasp, alongside the decoded hand motions. While the hyperbolic geodesic crosses empty regions in the latent space, the pullback geodesic adheres closely to the data support. As both geodesics start and end with the same grasps, their initial and final predictions match. However, the decoded hyperbolic geodesic shows high uncertainty in the middle part of the motion (see Figs. 5(a)-5(b) top right). This is due to the GP reverting to its non-informative mean in regions with no data, where the GP prior dominates the posterior, leading to maximum uncertainty. Consequently, the obtained hand motions display large deviations from the initial ring and final spherical grasps. In contrast, as shown in Fig. 5 and Table 2, the decoded pullback geodesic shows much lower uncertainty, as the geodesic stays closer to the data support. Note that this mirrors the behavior of the GPLVM Euclidean pullback geodesics compared to straight lines (see Fig. 8 in App. D.5 and Table 2), thus highlighting the successful adaptation of the pullback approach to hyperbolic spaces. Moreover, the hand motions obtained from the decoded pullback geodesics result in more realistic transitions that those obtained with the hyperbolic ones. Despite the low uncertainty, the predicted motions sometimes lack smoothness due to the fact that nearby latent points do not always correspond to equally similar hand configurations (i.e., similar joint values). This is mostly alleviated in the 3D latent space, which provides more room to accommodate the latent embeddings.